A small park is in the shape of a rectangle that measures 525 feet by 468 feet.
a. Find the number of feet of fencing that would be needed to enclose the park. Express your answer to the nearest foot.
b. If the entire park is to be painted with grass seed, find the number of square feet to be seeded. Express your answer to the correct number of significant digits based on the given dimensions.

Answer by checkley77(12844)

2*525+2*468=1,050+936=1,986 feet will be needed to enclose the park.
525*468=245,700 ft^2 is the area to be seeded.